find the equation of theline passing through the point (-1,-4) that is parallel to the line -9+5y=6.
I dont know whatto do with the equation without an x to make it into slope intercept form

Found 2 solutions by ewatrrr, ReadingBoosters:
Answer by ewatrrr(24785)


 
Hi,
find the equation of the line passing through the point (-1,-4)
that is parallel to the line -9+5y=6 0r 5y = 15, y = 3 (an horizontal line)
NEW line paralel to it passing through the point (-1,-4) would be y = -4
See Below:

Donations on my website are appreciated. Thanks. ewatrrr
http://math-is-magical.weebly.com/index.html

Answer by ReadingBoosters(3246)

-9+5y=6 as a linear equation
5y=15
y = 3, which is a straight horizontal line at y=3 passing though all x points.
There is no slope.
...
A parallel line would need to be also horizontal but passing (-1,-4), where the horizontal line passes along y=-4 at every x point, including -1.
Again, there is no slope; otherwise, they would eventually intersect.
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.